OpenStudy (anonymous):

what is 1=2/(3 cuberootx)

OpenStudy (anonymous):

x=(2/3)^3

OpenStudy (asnaseer):

is this your question?\[1=\frac{2}{3\sqrt[3]{x}}\]

jhonyy9 (jhonyy9):

so than 3cuberootx=2 cuberootx=2/3 x=8/27

OpenStudy (anonymous):

to asnaseer, yeah thats exactly how it is I just didn't know how to write it in

OpenStudy (asnaseer):

\[1=\frac{2}{3\sqrt[3]{x}}\]therefore:\[3\sqrt[3]{x}=2\]\[\sqrt[3]{x}=\frac{2}{3}\]so:\[x=(\frac{2}{3})^3=\frac{2^3}{3^3}=\frac{8}{27}\]as @cauchy and @jhonyy9 stated.